# Rule or Story, Which is a Better Commonsense Expression for Talking with Large Language Models?

### v1

@@ C.2 Error Analysis
C.3 Correlation Between Answer Accuracy and the Two Scores of Stories
We plot the correlation between answer accuracy and the commonsense score in Figure 9, and the BERT similarity score in Figure 10. The dashed lines are connected between the results of using stories and rules of the same datasets, which further examination reveals positive correlations of the two scores between using stories and using rules on most datasets
@@ E.4 An Example of Generated Stories Alongside the Self-SFT Iterations
Iteration
Story
Commonsense Score BERT Similarity Predicted Answer
0
Jane recalls a time when she was in her friend’s kitchen and saw her using a glue stick to attach a decoration to a cake. Jane watched as her friend carefully applied the glue to the back of the decoration, then pressed it onto the cake. Jane was impressed by how smoothly the glue stuck to the cake and how easily the decoration stayed in place. She had never thought about adults using glue sticks before, but now she could see how useful they could be for a variety of tasks.
0.657 0.827 E. kitchen drawer ✗
1
As a child, Jane used to spend hours crafting with glue sticks. She would often use them to attach paper, cardboard, and other materials to create various projects. She would usually keep a glue stick and a small container of glue near her workspace, which was usually a table in the living room or her bedroom. She would also sometimes use glue sticks to fix small broken objects around the house, such as toys or household items.
0.692 0.830 C. at school ✗
2
Jane was working as an intern at a local law firm. One day, her supervisor asked her to help organize the office supplies. As she opened the drawer of the desk she was assigned, Jane noticed a package of glue sticks neatly arranged next to the stapler and scissors. Curiosity piqued, she inquired about the glue sticks. Her supervisor explained that the office frequently used them for various tasks such as attaching documents to presentation boards or assembling reports. Jane was surprised but intrigued by this unconventional use of glue sticks in an office setting.
0.809 0.848 D. office ✓
3
I remember one specific experience where I saw adults using glue sticks. It was during my first day of an office internship. I was assigned to help organize some files in the storage room when I noticed a group of adults huddled around a table. Curiosity got the better of me, so I approached them to see what they were doing. To my surprise, they were using glue sticks to attach important documents onto presentation boards. It was fascinating to watch them carefully apply the glue in a controlled manner, ensuring the papers were securely fixed. I realized that adults indeed use glue sticks in an office setting for various tasks.
0.810 0.874 D. office ✓

### v2

@@ C.3 Error Analysis
C.4 Correlation Between Answer Accuracy and the Two Scores of Stories
We plot the correlation between answer accuracy and the commonsense score in Figure 9, and the BERT similarity score in Figure 10. The dashed lines are connected between the results of using stories and rules of the same datasets, which further reveal positive correlations of the two scores between using stories and using rules on most datasets
@@ C.4 Correlation Between Answer Accuracy and the Two Scores of Stories
C.5 Length of Stories and Rules
We calculate the average lengths of a single story and rule generated by the models used in our study. Lengths are calculated as the number of words while excluding stop words. The results are shown in Table 10. This analysis reveals that, on average, stories are longer than rules. However, this difference in length does not imply a difference in the amount of information conveyed. Stories tend to include longer sentences with detailed descriptions and personal feelings, incorporating adjectives and adverbs, while rules are typically more concise and straightforward. Despite these stylistic differences, both formats provide a comparable amount of commonsense information.
Model Story Length Rule Length
ChatGPT 82.6 65.5
Alpaca 50.1 40.7
Vicuna 62.2 49.9
Table 10: Average lengths (number of words) of generated stories and rules, excluding stop words.
@@ E.4 An Example of Generated Stories Alongside the Self-SFT Iterations
Iteration Story Commonsense Score BERT Similarity Predicted Answer
0 Jane recalls a time when she was in her friend’s kitchen and saw her using a glue stick to attach a decoration to a cake. Jane watched as her friend carefully applied the glue to the back of the decoration, then pressed it onto the cake. Jane was impressed by how smoothly the glue stuck to the cake and how easily the decoration stayed in place. She had never thought about adults using glue sticks before, but now she could see how useful they could be for a variety of tasks. 0.657 0.827 E. kitchen drawer ✗
1 As a child, Jane used to spend hours crafting with glue sticks. She would often use them to attach paper, cardboard, and other materials to create various projects. She would usually keep a glue stick and a small container of glue near her workspace, which was usually a table in the living room or her bedroom. She would also sometimes use glue sticks to fix small broken objects around the house, such as toys or household items. 0.692 0.830 C. at school ✗
2 Jane was working as an intern at a local law firm. One day, her supervisor asked her to help organize the office supplies. As she opened the drawer of the desk she was assigned, Jane noticed a package of glue sticks neatly arranged next to the stapler and scissors. Curiosity piqued, she inquired about the glue sticks. Her supervisor explained that the office frequently used them for various tasks such as attaching documents to presentation boards or assembling reports. Jane was surprised but intrigued by this unconventional use of glue sticks in an office setting. 0.809 0.848 D. office ✓
3 I remember one specific experience where I saw adults using glue sticks. It was during my first day of an office internship. I was assigned to help organize some files in the storage room when I noticed a group of adults huddled around a table. Curiosity got the better of me, so I approached them to see what they were doing. To my surprise, they were using glue sticks to attach important documents onto presentation boards. It was fascinating to watch them carefully apply the glue in a controlled manner, ensuring the papers were securely fixed. I realized that adults indeed use glue sticks in an office setting for various tasks. 0.810 0.874 D. office ✓